OpenStudy (anonymous):

7 3/4 - 3/8 = 7 1/4 Not Here 7 3/8 7 1/8 7 1/2

Directrix (directrix):

7 3/8 ----------- 31/4 - 3/8 = 62/8 - 3/8 = 59/8 = 7 and 3/8